# Local structural disorder in crystalline materials

### v1

@@ 5 Electron-phonon coupling
The anharmonic phonons obtained by accounting for local disorder, phonon self-energy corrections, or both can be coupled with the electronic structure of the locally disordered system, yielding the so called anharmonic electron-phonon coupling (Figure 3c). Electron-phonon coupling, in turn, affects both phonon and electron energies and lifetimes through self-energy corrections, as discussed extensively in Ref. 12. Figure 3d shows the effect of electron-phonon coupling on the band gap renormalization of cubic CsPbI3, as reported in Ref. 21. DFT calculations are performed within the ASDM using a $\times4$$\times$4 supercell. Results are averaged over 10 polymorphous configurations, weighted by their Boltzmann factors derived from free energies, thereby incorporating configurational entropy effects into the electron-phonon calculations. Employing polymorphous structures yields a positive band gap temperature coefficient, i.e., a band gap opening with increasing temperature. In contrast, calculations based on a monomorphous structure, not only underestimate the band gap, but also predict a band gap closing, due to an incomplete treatment of the Fan-Migdal self-energy contribution 19. In both cases, band gap renormalization is dominated by low energy optical modes (<<10 meV) with bending or rocking character, linked to B-X-B bond angle variations and octahedral tilting 19.
@@
Figure 4: Impact of local disorder on transport and polaron formation. (a) Calculated electron mobility of monomorphous and polymorphous cubic SrTiO3 compared with experiment. Reproduced with permission from Ref. 34, copyright (2024) by the American Physical Society. (b) Calculated average charge-carrier mobility of orthorhombic MAPbI3 and cubic CsPbI3 compated with experiment. Reproduced with permission from Ref. 35, copyright (2019) by the American Chemical Society. (c) Polaron charge density in CsPbBr3 calculated by PBE0 in a supercell after adding a negative charge. Figure adapted from Ref. 36 under $C$ BY-NC 4.0.
@@
Figure 4a compares calculated electron mobilities, obtained using monomorphous and polymorphous structures of SrTiO3, with experimental data as reported in Ref. 34. The mobilities are computed using perturbative methods based on the relaxation time approximation (RTA) and iterative solutions of the Boltzmann transport equation (IBTE), as implemented in the EPW 37 code. Anharmonic phonons are obtained via SSCHA. The effect of local disorder is incorporated by substituting the electronic eigenvalues computed for the monomorphous structure with those obtained from the polymorphous structure, while retaining the wavefunctions of the monomorphous configuration. Despite this approximation, local disorder reduces the electron mobility by 43% at 300K, significantly improving agreement with experiment. An even greater reduction is achieved when local disorder is combined with electron-phonon coupling effects on the electronic structure, lowering the initial mobility overestimation from 838% to 290% 34, a remarkable improvement. Zhou and Bernardi 38 using the Perturbo code 39 achieved excellent agreement with experimental electron mobilities of SrTiO3 by combining the retarded cumulant approach with the Kubo formula, capturing both coherent quasiparticles and incoherent polaronic effects through the spectral function. While their method treats dynamical electron-phonon interactions accurately, it neglects local structural disorder. Combining these two effects remains an open and promising direction for future research, particularly in understanding their joint impact on charge transport 40. Poncé et al 35 investigated the origin of low carrier mobilities in halide perovskites by performing EPW calculations within the RTA for the orthorhombic phase of MAPbI3 and cubic phase of CsPbBr3. Their results, presented in Figure 4b, illustrate the influence of local disorder on carrier mobilities in an indirect yet insightful manner. Employing the orthorhombic structure, which incorporates the effect of octahedral tilting similar to local disorder in the cubic phase, yields better agreement with experimental data 41, 42. This holds even for temperatures above 300 K, where the tetragonal to cubic phase transition takes place. Instead, using monomorphous cubic CsPbI3 leads to a significant increase of the calculated carrier mobilities, leading to the overestimation of the experimental data by a factror of 7.
@@
Figure 4c shows the calculated charge density of a large hole polaron in cubic CsPbBr3 using a $\times2$$\times$8 supercell. Polaron is a manifestation of electron-phonon coupling, where the carrier distorts the surrounding atomic structure and becomes dressed by phonons, forming a quasiparticle. The calculations were performed by removing one electron and performing a geometry relaxation using the hybrid PBE0 functional to explain large polarons observed experimentally using time-resolved optical Kerr effect spectroscopy 36. However, the computational approach employs a monomorphous cubic structure as the reference when introducing a charge and analyzing structural distortions. When a charged system is compared only to this idealized reference, the resulting lattice distortions may be overestimated or misattributed entirely to polaron formation. In reality, much of the distortion may already exist in the neutral, uncharged structure if modeled with a more physically accurate, symmetry-broken polymorphous configuration. Failing to use a locally disordered structure can lead to misleading conclusions about the nature, size, and stabilization energy of polarons in soft, anharmonic materials like CsPbBr3.
@@ 6 Future directions
The findings discussed in this Perspective establish the importance of incorporating local disorder through polymorphous structural models to accurately predict and understand the electronic, vibrational, and electron-phonon properties of soft, anharmonic high temperature phases of materials. Despite this progress, the theoretical treatment of electron-phonon coupling in the presence of positional polymorphism remains at an early stage, offering numerous opportunities for future research.
@@
One promising avenue is the integration of polymorphous supercells with state-of-the-art methodologies for electron-phonon coupling, such as the use of perturbative approaches implemented in state-of-the-art codes like EPW 37 and Perturbo 39. These methods typically assume high symmetry reference structures, limiting their predictive power in materials where symmetry breaking is intrinsic. Developing workflows with locally disordered configurations could significantly improve accuracy in modeling carrier lifetimes, charge transport, and phonon-assisted optical absorption. Regarding charge transport, another promising direction is to advance non-perturbative supercell approaches for anharmonic electron-phonon coupling, allowing the evaluation of carrier mobilities that include temperature-dependent band gap renormalization on top of local disorder effects 40. An advantage of supercell approaches, despite lacking the elegance of fine Brillouin zone sampling, is their ability to seamlessly treat anharmonicity and electron-phonon coupling on the same footing, without relying on PES derivatives along harmonic phonons 38.
@@
Another exciting direction lies in exploring non-equilibrium carrier dynamics in disordered lattices. First-principles simulations of carrier-phonon and phonon-phonon interactions in the ultrafast regime 43, 44, can be extended to polymorphous structures to probe transient phenomena such as hot carrier relaxation, carrier localization, and coherent phonon generation, as well as to uncover scattering channels that are normally forbidden by symmetry in the monomorphous case. In parallel, excitonic effects in materials with strong local disorder remain largely unexplored; investigating how positional polymorphism influences exciton binding energies, lifetimes, and dissociation dynamics could offer new insights into light-matter interactions and optical response. Similarly, Auger recombination and radiative carrier lifetimes, which are central to optoelectronic device performance 45, can be re-evaluated.
@@
The application of these frameworks to a broader class of materials, including layered/2D perovskites, double perovskites, antiperovskites, and quantum dots is another frontier. These systems often feature pronounced anisotropy and molecular complexity, making them ideal candidates for a polymorphous description. Particularly, the interplay of molecular cation dynamics with inorganic octahedral tilting and quantum confinement, deserves detailed investigation, potentially requiring the development of hybrid methodologies capable of treating the different sources of local disorder consistently 21.
@@
It will also be interesting to compile databases of known and predicted anharmonic materials and conducting high-throughput first-principles studies to systematically explore their polymorphous structural networks across a broad chemical space 18. This effort would lay the groundwork for identifying trends and guiding the discovery of materials where local disorder plays a functional role. Moreover, given the large configurational space associated with polymorphous networks, machine learning force fields offer a powerful tool to accelerate structure exploration in large supercells and thus capture long range disorder as well as perform classification and even optoelectronic property prediction. This effort could also provide insight into how polar symmetry breaking is entangled with local disorder in complex material systems, including but not limited to Ge-based perovskites.
@@
It is worth noting that the polymorphous framework has certain limitations. These include the lack of an explicit treatment of relaxation processes detectable by low-energy spectroscopies and the restricted supercell sizes, which may underestimate correlation lengths and produce simulated Bragg-like features instead of experimental diffuse scattering.
@@
The full integration of the SCP theory with polymorphous structures offers a rigorous path toward modeling strongly anharmonic vibrations and demonstrate further the breakdown of the phonon quasiparticle picture. Such a framework would enable temperature-dependent phonon dispersions, lifetimes, thermal conductivity, and electron-phonon coupling coefficients to be computed on disordered environments, resolving current ambiguities in interpreting temperature-induced property changes. Modeling thermal conductivity in locally disordered materials is critical, since the mean free paths can be reduced down to interatomic distances 9, and scattering becomes dominated by diffuson-like modes. This opens a path toward predicting ultralow thermal conductivity in materials exhibiting entropy-driven structural complexity and glass-like behavior, despite their crystalline nature. Thermal conductivity relies on higher-order interatomic force constants, particularly third- and fourth-order terms governing phonon-phonon scattering, making machine learning force fields an essential tool for accelerating their computation in complex, disordered systems. Furthermore, extending local disorder to consider temperature-driven phase transitions and disorder-enabled functionalities such as ferroelectricity, spin polymorphism in paramagnetic phases, ionic transport, or polaron formation represents a rich area for computational development.
@@
From the experimental perspective, advancing characterization methods that are sensitive to local disorder is equally essential. While PDF analysis remains a powerful tool to probe local structure, complementary techniques such as high-temperature Raman spectroscopy, nuclear magnetic resonance measurements, inelastic neutron scattering for phonon DOS, and diffuse scattering can offer critical insights into vibrational dynamics and structural fluctuations in disordered systems. These methods provide additional observables that can be directly compared with first-principles predictions, allowing for more rigorous validation of polymorphous models, as well as elucidate long-range correlated disorder. Establishing such cross-validation pipelines between theory and experiment is key to uncovering the main mechanisms underpinning the efficiency of functional materials.

### v2

@@ 4 Electron-phonon coupling
The anharmonic phonons obtained by accounting for local disorder, phonon self-energy corrections, or both can be coupled with the electronic structure of the locally disordered system, yielding the so called anharmonic electron-phonon coupling (Figure 3c). Electron-phonon coupling, in turn, affects both phonon and electron energies and lifetimes through self-energy corrections, as discussed extensively in Ref. 15. Figure 3d shows the effect of electron-phonon coupling on the band gap renormalization of cubic CsPbI3, as reported in Ref. 24. DFT calculations are performed within the ASDM using a $\times4$$\times$4 supercell. Results are averaged over 10 polymorphous configurations, weighted by their Boltzmann factors derived from free energies, thereby incorporating configurational entropy effects into the electron-phonon calculations. Employing polymorphous structures yields a positive band gap temperature coefficient, i.e., a band gap opening with increasing temperature. In contrast, calculations based on a monomorphous structure, not only underestimate the band gap, but also predict a band gap closing, due to an incomplete treatment of the Fan-Migdal self-energy contribution 22. In both cases, band gap renormalization is dominated by low energy optical modes (<<10 meV) with bending or rocking character, linked to B-X-B bond angle variations and octahedral tilting 22.
@@
Figure 4: Impact of local disorder on transport and polaron formation. (a) Calculated electron mobility of monomorphous and polymorphous cubic SrTiO3 compared with experiment. Reproduced with permission from Ref. 43, copyright (2024) by the American Physical Society. (b) Calculated average charge-carrier mobility of orthorhombic MAPbI3 and cubic CsPbI3 compated with experiment. Reproduced with permission from Ref. 44, copyright (2019) by the American Chemical Society. (c) Polaron charge density in CsPbBr3 calculated by PBE0 in a supercell after adding a negative charge. Figure adapted from Ref. 45, copyright (2017) by the American Association for the Advancement of Science.
@@
Figure 4a compares calculated electron mobilities, obtained using monomorphous and polymorphous structures of SrTiO3, with experimental data as reported in Ref. 43. The mobilities are computed using perturbative methods based on the relaxation time approximation (RTA) and iterative solutions of the Boltzmann transport equation (IBTE), as implemented in the EPW 46 code. Anharmonic phonons are obtained via SSCHA. The effect of local disorder is incorporated by substituting the electronic eigenvalues computed for the monomorphous structure with those obtained from the polymorphous structure, while retaining the wavefunctions of the monomorphous configuration. Despite this approximation, local disorder reduces the electron mobility by 43% at 300K, significantly improving agreement with experiment. An even greater reduction is achieved when local disorder is combined with electron-phonon coupling effects on the electronic structure, lowering the initial mobility overestimation from 838% to 290% 43, a remarkable improvement. Zhou and Bernardi 47 using the Perturbo code 48 achieved excellent agreement with experimental electron mobilities of SrTiO3 by combining the retarded cumulant approach with the Kubo formula, capturing both coherent quasiparticles and incoherent polaronic effects through the spectral function. While their method treats dynamical electron-phonon interactions accurately, it neglects local structural disorder. Combining these two effects remains an open and promising direction for future research, particularly in understanding their joint impact on charge transport 49. Poncé et al 44 investigated the origin of low carrier mobilities in halide perovskites by performing EPW calculations within the RTA for the orthorhombic phase of MAPbI3 and cubic phase of CsPbBr3. Their results, presented in Figure 4b, illustrate the influence of local disorder on carrier mobilities in an indirect yet insightful manner. Employing the orthorhombic structure, which incorporates the effect of octahedral tilting similar to local disorder in the cubic phase, yields better agreement with experimental data 50, 51. This holds even for temperatures above 300 K, where the tetragonal to cubic phase transition takes place. Instead, using monomorphous cubic CsPbI3 leads to a significant increase of the calculated carrier mobilities, leading to the overestimation of the experimental data by a factror of 7.
@@
Figure 4c shows the calculated charge density of a large hole polaron in cubic CsPbBr3 using a $\times2$$\times$8 supercell. Polaron is a manifestation of electron-phonon coupling, where the carrier distorts the surrounding atomic structure and becomes dressed by phonons, forming a quasiparticle. The calculations were performed by removing one electron and performing a geometry relaxation using the hybrid PBE0 functional to explain large polarons observed experimentally using time-resolved optical Kerr effect spectroscopy 45. However, the computational approach employs a monomorphous cubic structure as the reference when introducing a charge and analyzing structural distortions. When a charged system is compared only to an idealized reference, the resulting lattice distortions may be overestimated or misattributed entirely to polaron formation. In reality, much of the distortion may already exist in the neutral, uncharged structure if modeled with a more physically accurate, symmetry-broken polymorphous configuration. Failing to use a locally disordered structure can lead to misleading conclusions about the nature, size, and stabilization energy of polarons in soft, anharmonic materials like CsPbBr3.
@@ 5 Future directions
The findings discussed in this Perspective establish the importance of incorporating local disorder through polymorphous structural models to accurately predict and understand the electronic, vibrational, and electron-phonon properties of soft, anharmonic high temperature phases of materials. Despite the progress, the theoretical treatment of electron-phonon coupling in the presence of positional polymorphism remains at an early stage, offering numerous opportunities for future research.
@@
One promising avenue is the integration of polymorphous supercells with state-of-the-art methodologies for electron-phonon coupling, such as the use of perturbative approaches implemented in state-of-the-art codes like EPW 46 and Perturbo 48. These methods typically assume high symmetry reference structures, limiting their predictive power in materials where symmetry breaking is intrinsic. Developing workflows with locally disordered configurations could significantly improve accuracy in modeling carrier lifetimes, charge transport, and phonon-assisted optical absorption. Regarding charge transport, another promising direction is to advance non-perturbative supercell approaches for anharmonic electron-phonon coupling, allowing the evaluation of carrier mobilities that include temperature-dependent band gap renormalization on top of local disorder effects 49. An advantage of supercell approaches, despite lacking the elegance of fine Brillouin zone sampling, is their ability to seamlessly treat anharmonicity and electron-phonon coupling on the same footing, without relying on PES derivatives along harmonic phonons 47.
@@
Another exciting direction lies in exploring non-equilibrium carrier dynamics in disordered lattices. First-principles simulations of carrier-phonon and phonon-phonon interactions in the ultrafast regime 52, can be extended to polymorphous structures to probe transient phenomena such as hot carrier relaxation, carrier localization, and coherent phonon generation, as well as to uncover scattering channels that are normally forbidden by symmetry in the monomorphous case. In parallel, excitonic effects in materials with strong local disorder remain largely unexplored; investigating how positional polymorphism influences exciton binding energies, lifetimes, and dissociation dynamics could offer new insights into light-matter interactions and optical response. Similarly, Auger recombination and radiative carrier lifetimes, which are central to optoelectronic device performance 53, can be re-evaluated.
@@
The application of these frameworks to a broader class of materials, including layered/2D perovskites, double perovskites, antiperovskites, and quantum dots is another frontier. These systems often feature pronounced anisotropy and molecular complexity, making them ideal candidates for a polymorphous description. Particularly, the interplay of molecular cation dynamics with inorganic octahedral tilting and quantum confinement, deserves detailed investigation, potentially requiring the development of hybrid methodologies capable of treating the different sources of local disorder consistently 24.
@@
It will also be interesting to compile databases of known and predicted anharmonic materials and conducting high-throughput first-principles studies to systematically explore their polymorphous structural networks across a broad chemical space 21. This effort would lay the groundwork for identifying trends and guiding the discovery of materials where local disorder plays a functional role. Moreover, given the large configurational space associated with polymorphous networks, machine learning force fields offer a powerful tool to accelerate structure exploration in large supercells and thus capture long range disorder as well as perform classification and even optoelectronic property prediction. This effort could also provide insight into how polar symmetry breaking is entangled with local disorder in complex material systems, including but not limited to Ge-based perovskites.
@@
It is worth noting that the polymorphous framework has certain limitations. These include the lack of an explicit treatment of relaxation processes detectable by low-energy spectroscopies and the restricted supercell sizes, which may underestimate correlation lengths and produce simulated Bragg-like features instead of experimental diffuse scattering. Furthermore, in its current implementation, local disorder is explored at fixed lattice parameters, neglecting strain-structure coupling and potentially biasing the accessible minima. Future extensions should incorporate variable-cell relaxations and larger supercells. Another limitation of the polymorphous framework is its quasi-static, Born-Oppenheimer treatment, which neglects non-adiabatic forces which can govern photoinduced structural dynamics 54.
@@
The full integration of the SCP theory with polymorphous structures offers a rigorous path toward modeling strongly anharmonic vibrations and demonstrate further the breakdown of the phonon quasiparticle picture. Such a framework would enable temperature-dependent phonon dispersions, lifetimes, thermal conductivity, and electron-phonon coupling coefficients to be computed on disordered environments, resolving current ambiguities in interpreting temperature-induced property changes. Modeling thermal conductivity in locally disordered materials is critical, since the mean free paths can be reduced down to interatomic distances 10, and scattering becomes dominated by diffuson-like modes. This opens a path toward predicting ultralow thermal conductivity in materials exhibiting entropy-driven structural complexity and glass-like behavior, despite their crystalline nature 38, 36, 19. Thermal conductivity relies on higher-order interatomic force constants, particularly third- and fourth-order terms governing phonon-phonon scattering, making machine learning force fields an essential tool for accelerating their computation in complex, disordered systems. Furthermore, extending local disorder to consider temperature-driven phase transitions and disorder-enabled functionalities such as ferroelectricity, spin polymorphism in paramagnetic phases, ionic transport, or polaron formation represents a rich area for computational development.
@@
From the experimental perspective, advancing characterization methods that are sensitive to local disorder is equally essential. While PDF analysis remains a powerful tool to probe local structure, complementary techniques such as high-temperature Raman spectroscopy, nuclear magnetic resonance measurements, inelastic neutron scattering for phonon DOS, and diffuse scattering can offer critical insights into vibrational dynamics and structural fluctuations in disordered systems 6, 12, 11, 37. These methods provide additional observables that can be directly compared with first-principles predictions, allowing for more rigorous validation of polymorphous models, as well as elucidate long-range correlated disorder. Establishing such cross-validation pipelines between theory and experiment is key to uncovering the main mechanisms underpinning the efficiency of functional materials.
6 Acknowledgments
We acknowledge computational resources from the EuroHPC Joint Undertaking and supercomputer LUMI [https://lumi-supercomputer.eu/], hosted by CSC (Finland) and the LUMI consortium through a EuroHPC Extreme Scale Access call.